OpenStudy (anonymous):

simplify the radical expression. 5. √250h^4k^5 6. √14q x 2√4q 7. √63x^15y^9/7xy^11 8. √3+4√3 5√6 5√3 3√6 3√3 9. 2√6+3√96 14√6 14√96 5√96 50√6 10. (8 +√11 )(8 –√11 ) 53 75 + 16√11 –57 64 + √11

OpenStudy (anonymous):

\[\sqrt{250h ^{4}k^{5}} = \sqrt{(25)(10)h^{4}k^{5}} = 5h^{2}k^{2}\sqrt{10k}\]

OpenStudy (anonymous):

\[\sqrt{14q}. 2 .\sqrt{4q} = 2.\sqrt{(14)(4)q^{2}} = (2)(2)(q)\sqrt{14} = 4q \sqrt{14}\]